# How digital will the future be? Analysis of prospective scenarios

### v1

@@ 4   Results
a) Digital technologies by scenario b) Digital technologies by studies
Refer to caption Refer to caption
c) Applications by scenario d) Applications by studies
Refer to caption Refer to caption
e) Infrastructures by scenario f) Infrastructures by studies
Figure 2: Summary of our scenario analysis for ICT. a),c),e): Number of scenarios among the 35 explicitly mentioning an ICT variable. b),d),f): Number of prospective studies among the 14 explicitly mentioning an ICT variable. For complete details, we refer the readers to our complementary material.
@@
4.1 Digital technologies and ICT characteristics
@@
As a consequence of the use of data by digital technologies, 5 scenarios mention open data in four studies: (in studies [Ademe, 2021, Arup, 2019, European Commission, 2019, Creutzig et al., 2022]). Open data is essential in the ”Deliberate for the good” scenario of the ICT-specific study [Creutzig et al., 2022]. It comes with the idea that sustainable ICT requires more equity and collective data management.
Concerning the environmental impacts of digital technologies, four scenarios specifically mention mandatory life-cycle assessments of all new services and equipment (in studies [Arup, 2019, The Shift Project, 2020, European Commission, 2019, Deron and McDonald, 2022]).
4.2 Applications of digital technologies
@@ 4.2.1 Main applications
4.2.2 Applications only discussed in a few scenarios
@@
  • e-commerce: We only saw e-commerce in four studies ([The Shift Project, 2020, Ademe, 2021, Danish Design Center, 2020]) with an increase (in particular for business-as-usual scenarios) or a decrease (for more sober scenarios). Supply chain management is mentioned in [IPCC, 2022] but not specifically for e-commerce.
  • Information on climate change: ICT is important in some scenarios to raise awareness of climate issues and guide citizens. Digital media can be used to promote a culture of sustainability [Ministère de la Transition écologique et solidaire, 2020, IPCC, 2022, European Commission, 2019] and to inform about one’s own emissions and carbon quota [European Environment Agency, 2021, Arup, 2019].
  • Social media: ICT is also still used for communication and social media and explicitly mentioned as such in the Regional Cooperations scenario of [Ademe, 2021], in the Ecotopia and Technocracy scenarios of [European Environment Agency, 2021], and in [CNIL, 2021]. In these scenarios, digital technologies enable citizens to interact continuously.
  • Leisure: Leisure relying on digital technologies is discussed in several studies, with different visions in different scenarios. In [Ademe, 2021], the frugal scenario assumes less digital leisure, while the Restauration Gamble one assumes more and more digital leisure. Conversely, the [négaWatt, 2021] scenario, which is a relatively sober scenario, supposes that digitalization continues for leisure activities. Even more surprising, the Ecotopia scenario of the study [European Environment Agency, 2021] scenario, which is, in essence, closer to nature, considers an increased presence of the digital sphere.
    Digital technologies can also be used to improve the living environment, as in the Humans Inc. scenario of [Arup, 2019], which mentions digital walls displaying images of a clean waterfront to protect residents from unpleasant views.
  • Surveillance: In general studies, surveillance was only found in the Extinction express scenario of [Arup, 2019] with the use of micro-drones by the police and in centralized scenarios from [Danish Design Center, 2020], with AI or autonomous units. Naturally, it is discussed more deeply in [CNIL, 2021], the main topic being data protection and individual privacy. As such, data surveillance is a central variable of this study.
@@
4.2.3 Sobriety in digital technologies
Some scenarios call for sobriety and therefore, bring sobriety within the digital sector inherently. This is visible, for instance, in the sufficiency scenario of [RTE, 2022] with less digital leisure and less digital advertising. In the Frugal Generation scenario from [Ademe, 2021], digital sobriety implies no new data center development and a stop of data growth. Another aspect of this sobriety concerns equipment repairability by users, which is discussed in many scenarios to increase equipment lifetime. In [Deron and McDonald, 2022], a reduction of the number of pieces of equipment is also possible with the development of equipment sharing in companies.
Some sober scenarios [négaWatt, 2021, The Shift Project, 2020] consider digital sobriety from a digital technologies usage angle. The proposal is to control usage and limit data flows. In the scenario from [The Shift Project, 2020], digital sobriety includes reducing video resolutions or cloud gaming, or bounding equipment rate per person. But, on the contrary, this same scenario assumes an increase in the use of ICT in many sectors (such as higher education or defense). [Creutzig et al., 2022]propose an extensive application of ICT to match sustainable development goals.
5   Discussions
The first observation of our survey was that there is no scenario little or no ICT. Due to the increasing environmental impacts of digital technologies [Freitag et al., 2021], it is essential to think about their place in an environmentally-constrained world. The role of digitalization in the anthropocene was also the starting point of [Creutzig et al., 2022]. In addition, ICT is related to significant physical, resilience, and environmental challenges that should drive its future developments.
In the following, we discuss how the studies considered these challenges and what is currently missing when considering ICT in prospective studies.
@@ 5   Discussions
All scenarios admit a digital world quite similar to that of today. It can be more minimalistic or developed, sometimes with sobriety and efficiency, but no significant change exists. Many scenarios give a central place to IT for Green applications (smart farming, smart building, ecological monitoring). Some structural changes come from innovations in ICT, such as teleworking or autonomous driving.
As our goal was to study the place of digital technologies in future scenarios, we defined general ICT-specific variables to analyze if some scenarios consider them as the starting point. These variables encompass the growth of data, location of data centers, or sobriety usage. ICT variables (section 3.1) can be seen as parameters that can be changed to achieve different scenarios. They could indeed serve as central hypotheses to design general scenarios, i.e. being one of the causes in the scenario. An example is the following: because there will be much efficiency in digital technologies in the future, these technologies will be more deployed. Or, because of a limitation on the number of data centers, digital technologies have stopped being applied to some sectors of society. However, no general study has built upon these considerations. ICT variables remain only consequences of other aspects of these scenarios. While ICT-centered studies [Creutzig et al., 2022, CNIL, 2021, Deron and McDonald, 2022] build upon these ICT variables, their scenarios hardly mention the application of ICT in various areas of society.
Scenarios rarely look at ICT with a comprehensive perspective. The studies often provide a citizen’s point of view on ICT more than companies, organizations, or society one. Not having a more systemic view of society implies that ICT is very application-oriented, omitting all the interconnection with infrastructures and technologies. For example, [The Shift Project, 2020] offers a specific plan for the future of digital technologies, but separated from other sectors of activities and with few interactions. While digital sobriety is encouraged, ICT’s place in other sectors is hardly discussed, even when it is already deployed.
@@
In addition, ICT is related to significant physical, resilience, and environmental challenges: ICT infrastructures, for example, are very vulnerable to extreme climate events. This vulnerability is only very rarely evoked in the scenarios. Only [The Shift Project, 2020] states that ”The ubiquity of digital technology makes its resilience crucial to many aspects of society”, and this variable is absent from all scenarios. Not considering the interaction of digital technologies with other sectors of activities may lead to inconsistent scenarios. Furthermore, digital technologies are often seen as crucial for decarbonization. Its central place in some decarbonization strategies with regard to its vulnerability is never discussed.
Of course, it is difficult for a prospective study to cover all areas of society. Nevertheless, those that serve essential needs such as drinking, eating, housing, or health care might be the ones that are the most important to develop imaginaries and make the right decisions. Many of these essential areas already depend on ICT today. But no study mentions them all. We can think for example, about all supply chains or water supply and management. We can, therefore, only guess that these areas may also follow business-as-usual or sobriety tendencies. This lack of details raises questions regarding the transformations that an increase or decrease in digital technologies may imply for these applications and people working in these fields. No study questions the resilience of society towards these digital technologies.
Another aspect that is rarely tackled in scenarios is cybersecurity. Another aspect that scenarios rarely tackle is cybersecurity. The capacity of the technologies to resist cyber-attacks is essential, especially in scenarios with more fragmentation of the world.
@@
Developing ICT in many sectors of activities, as proposed in some scenarios, would likely necessitate the installation of a significant amount of new digital infrastructures. Only some scenarios mention the availability or criticity of resources, pure water, energy, or land (e.g., with resource colonies on the moon and deep sea mining in Extinction express of [Arup, 2019], or in [RTE, 2022] scenarios for energy production). This physical reality behind digital technologies is most often forgotten.
Relying on ICT in the future requires looking at energy and resources supply, which geopolitical tensions, collaborations between countries, or changes in governance may hinder. These tensions are discussed in French National studies (e.g. RTE and negaWatt) and in the narrative study Arup, but do not seem to influence their integration of ICT. The Extinction express scenario of the latter describes resource colonies on the moon and in the deep sea.
To the best of our knowledge, no study, however, mentions the geopolitical aspects related to the location of data centers and the installation of ICT infrastructures (e.g., underwater cables or satellites).
@@
Surprisingly, many technologies that have started to receive attention and funding today are not mentioned in any of the studies, which build upon existing technologies only. For example, there is very little about 6G, and nothing on quantum informatics, intermittent computing, or DNA data storage. It is difficult to imagine how these innovations in ICT could be integrated in the scenarios. An example of studies where disruptive technologies (e.g., cyborg botany or olfactory notifications) are considered in scenarios has been presented by French designers333 https://b-com.com/actualite/immersion-dans-le-numerique-responsable-de-2040. Note that we did not include this work in our analysis because of a lack of details on the methodology. Significant changes in the use of technology, by using low-tech solutions for example, are also mostly absent from scenarios.
In [Erdmann and Hilty, 2010], the authors argued that forthcoming macroeconomic studies on ICT should account for first-, second-, and third-order effects, reflect error margins resulting from data uncertainty and use scenario techniques to explore future uncertainty. However, the studies we analyzed generally do not, even when ICT-centered. Some studies point to possible rebound effects of digital technologies, but indirect effects of ICT and their assessment are never included in the scenarios.
@@
In continuity with the previous point, very few general scenarios are peripheral, i.e., include unlikely and extreme events and consider a discontinuous path to the future. Generally, disruptive scenarios are narratives. When looking specifically at ICT in general scenarios, it is pretty similar to today’s. Our relationship to digital technologies is not deeply questioned, which may be why ICT is present in all scenarios without breakthrough innovations.
@@
The studies were all made from a climate change perspective, omitting biodiversity losses and other planetary boundaries and focusing on human futures only without considering other species. These environmental problems may have other consequences on future scenarios. One that is directly related to ICT is water scarcity. Indeed, the manufacturing of ICT equipment and data center cooling have a substantial ecological footprint in terms of water use. Prospective studies on ICT should, therefore, broaden their narrations.
@@
Most innovations in the ICT sector and major industries of the domain are from Global North countries. As a consequence, the studies we have chosen were made in OECD countries, and this clearly has an influence on scenarios. Some studies (e.g., Arup) provide visions within the world and have been designed with people from several countries worldwide. Nevertheless, our work is more likely not representative of visions from Global South studies, and future works should consider how to imagine ICT for the future in Global South countries.
6   Conclusion
In this paper, we have analyzed the place of ICT in prospective scenarios. We have highlighted some variables that influence the development of digital technologies and the areas of applications where ICT might be essential in the future. Our analysis demonstrates that the feasibility and consequences of the increasing development of digital technologies are not sufficiently considered. Through our discussions, we have shown that the studies hardly question our relationships to digital technologies or the applications of ICT in the context of the Anthropocene. This shortcoming limits the possibility of imagining more diverse plausible futures for the ICT sector. There is, therefore, a need to design prospective studies that could offer this more diverse and systemic view of ICT and highlight which undone science [Hess, 2015] should receive more attention. Prospective studies with a sustainability perspective for ICT can i) offer a more diverse and systemic view of the future of digital technologies; ii) highlight the need for discussing, structuring, or funding. As an example, [Deron and McDonald, 2022] indeed demonstrate that, from prospective co-design workshops, it is possible to find milestones to reach a desirable future of digital technologies. The milestones show samples of undone sciences [Hess, 2015] for which social and industrial movements, funding, and research are needed for a desirable and sustainable ICT.
Appendix A Summary of studies reviewed for this paper
Table 2: Summary of studies reviewed for this paper
Study
Summary & Scenarios
IPCC 2022
Report from IPCC Group III. This document does not correspond to a unique scenario, but describes the mitigation options.

### v3

@@ 4. Results for ICT variables
a) Infrastructures by scenario b) Digital technologies by scenario
c) Applications by scenario
Figure 2. Summary of our scenario analysis for ICT. Each diagram indicates the number of scenarios among the 35 explicitly mentioning an ICT variable. For complete details, we refer the readers to our complementary material.
@@
4.1. Digital technologies and ICT characteristics
@@
As a consequence of the use of data by digital technologies, 5 scenarios mention open data in four studies: (in studies (Ademe, 2021; Arup, 2019; European Commission, 2019; Creutzig et al., 2022)). Open data is essential in the ”Deliberate for the good” scenario of the ICT-specific study (Creutzig et al., 2022). It comes with the idea that sustainable ICT requires more equity and collective data management.
Concerning the environmental impacts of digital technologies, four scenarios specifically mention mandatory life-cycle assessments of all new services and equipment (in studies (Arup, 2019; Shift Project, 2020; European Commission, 2019; Deron and McDonald, 2022)).
4.2. Applications of digital technologies
@@ 4.2.1. Main applications
4.2.2. Applications only discussed in a few scenarios
@@
  • e-commerce: We only saw e-commerce in four studies ((Shift Project, 2020; Ademe, 2021; Danish Design Center, 2020)) with an increase (in particular for business-as-usual scenarios) or a decrease (for more sober scenarios). Supply chain management is mentioned in (IPCC, 2022) but not specifically for e-commerce.
  • Information on climate change: ICT is important in some scenarios to raise awareness of climate issues and guide citizens. Digital media can be used to promote a culture of sustainability (Ministère de la Transition écologique et solidaire, 2020; IPCC, 2022; European Commission, 2019) and to inform about one’s own emissions and carbon quota (European Environment Agency, 2021; Arup, 2019).
  • Social media: ICT is also still used for communication and social media and explicitly mentioned as such in the Regional Cooperations scenario of (Ademe, 2021), in the Ecotopia and Technocracy scenarios of (European Environment Agency, 2021), and in (CNIL, 2021). In these scenarios, digital technologies enable citizens to interact continuously.
  • Leisure: Leisure relying on digital technologies is discussed in several studies, with different visions in different scenarios. In (Ademe, 2021), the frugal scenario assumes less digital leisure, while the Restauration Gamble one assumes more and more digital leisure. Conversely, the (négaWatt, 2021) scenario, which is a relatively sober scenario, supposes that digitalization continues for leisure activities. Even more surprising, the Ecotopia scenario of the study (European Environment Agency, 2021), which is, in essence, closer to nature, considers an increased presence of the digital sphere.
    Digital technologies can also be used to improve the living environment, as in the Humans Inc. scenario of (Arup, 2019), which mentions digital walls displaying images of a clean waterfront to protect residents from unpleasant views.
  • Surveillance: In general studies, surveillance was only found in the Extinction express scenario of (Arup, 2019) with the use of micro-drones by the police and in centralized scenarios from (Danish Design Center, 2020), with AI or autonomous units. Naturally, it is discussed more deeply in (CNIL, 2021), the main topic being data protection and individual privacy. As such, data surveillance is a central variable of this study.
@@
4.2.3. Sobriety in digital technologies
Some scenarios call for sobriety and therefore, bring sobriety within the digital sector inherently. This is visible, for instance, in the sufficiency scenario of (RTE, 2022) with less digital leisure and less digital advertising. In the Frugal Generation scenario from (Ademe, 2021), digital sobriety implies no new data center development and a stop of data growth. Another aspect of this sobriety concerns equipment repairability by users, which is discussed in many scenarios to increase equipment lifetime. In (Deron and McDonald, 2022), a reduction of the number of pieces of equipment is also possible with the development of equipment sharing in companies.
Some sober scenarios (négaWatt, 2021; Shift Project, 2020) consider digital sobriety from a digital technologies usage angle. The proposal is to control usage and limit data flows. In the scenario from (Shift Project, 2020), digital sobriety includes reducing video resolutions or cloud gaming, or bounding equipment rate per person. But, on the contrary, this same scenario assumes an increase in the use of ICT in many sectors (such as higher education or defense). (Creutzig et al., 2022) propose an extensive application of ICT to match sustainable development goals. In (Deron and McDonald, 2022), business models around ICT are questioned to achieve more sobriety. A milestone specifically proposes to limit the influence of technology giants and to regulate the attention economy model.
5. Discussion
In the previous sections, we analyzed the kind of ICT included in the scenarios according to the typology of variables we defined. This section gives more general feedback on the scenarios, particularly on what we considered missing concerning the place of ICT.
A first general finding is that every scenario involves ICT, but without discussing the challenges of using ICT in an environmentally-constrained world. However ICT has a growing environmental footprint (Freitag et al., 2021), and raises resilience issues that should be taken into account. This situation calls for the role of digitalization in the context of anthropocene to be reexamined, as (Creutzig et al., 2022) advocates.
In the following discussion, we examine how the studies considered these challenges and what is currently missing concerning ICT in prospective studies.
@@ 5. Discussion
All scenarios admit a digital world quite similar to that of today. Although the imagined worlds can be more minimalistic or developed, sometimes with sobriety and efficiency, no significant change exists in their equipment and applications of IT. Some structural changes come from innovations in ICT, such as teleworking or autonomous driving.
As our goal was to study the place of digital technologies in future scenarios, we defined general ICT-specific variables to analyze if some scenarios consider them as the starting point. These variables encompass the growth of data, location of data centers, or sobriety usage. ICT variables (section 3.2) can be seen as parameters that can be changed to achieve different scenarios. They may serve as central hypotheses to design general scenarios, i.e., being one of the causes in the scenario. Two possible examples include 1) because there will be much efficiency in digital technologies in the future, these technologies will be more deployed; and 2) because of a limitation on the number of data centers, digital technologies have stopped being applied to some sectors of society. However, no general study has built upon these considerations. ICT variables remain only consequences of other aspects of these scenarios. While ICT-centered studies (Creutzig et al., 2022; CNIL, 2021; Deron and McDonald, 2022) build upon these ICT variables, their scenarios hardly mention the application of ICT in various areas of society. Authors in (Eriksson, 2018) advise to reconsider how ICT affect not only direct users, but other stakeholders, the biosphere, and coming generations. This is lacking in most scenarios we have studied, but several scenarios propose to regulate usages, data economy, attention economy (e.g. (Creutzig et al., 2022; Shift Project, 2020; Deron and McDonald, 2022)).
Scenarios rarely look at ICT with a comprehensive perspective, instead providing more often a citizen’s point of view on ICT more than companies, organizations, or society one. Not having a more systemic view of society implies that ICT is very application-oriented, omitting all the interconnection with infrastructures and technologies. For example, (Shift Project, 2020) offers a specific plan for the future of digital technologies, but separated from other sectors of activities and with few interactions. While digital sobriety is encouraged, ICT’s place in other sectors is hardly discussed, even when it is already deployed.
In (Erdmann and Hilty, 2010), the authors argued that forthcoming macroeconomic studies on ICT should account for first-, second-, and third-order effects, reflect error margins resulting from data uncertainty and use scenario techniques to explore future uncertainty. However, the studies we analyzed generally do not, even when ICT-centered. Some studies point to possible rebound effects of digital technologies, but indirect effects of ICT and their assessment are never included in the scenarios.
@@
Many sectors that cover essential needs such as drinking, eating, housing, or health care depend on ICT today, and imagining their future is thus important to make the right decisions. No study covers them all concurrently yet, and for example supply chains or water supply management are never detailed. This lack of details raises questions regarding the transformations that an increase or decrease in digital technologies may imply for these applications and people working in these fields.
In addition, ICT can be impacted by significant physical, resilience, and environmental challenges: ICT infrastructures, for example, are very vulnerable to extreme climate events. This vulnerability is very rarely evoked in the scenarios, and only (Shift Project, 2020) has stated that ”The ubiquity of digital technology makes its resilience crucial to many aspects of society”. As digital technologies are often seen as crucial for decarbonization of other sectors, not considering their vulnerability may lead to inconsistent scenarios. The absence of a resilience variable thus constitutes a limitation of current studies.
Cybersecurity is another aspect that is rarely tackled in scenarios. The capacity of technologies to resist cyber-attacks is essential, especially in scenarios with more fragmentation of the world.
@@
Developing ICT in many sectors of activities, as proposed in some scenarios, would likely necessitate the installation of a significant amount of new digital infrastructures. Only some scenarios mention the availability of critical resources, pure water, energy, or land (e.g., with resource colonies on the Moon and deep sea mining in Extinction express of (Arup, 2019), or in (RTE, 2022) scenarios for energy production). This physical reality behind digital technologies is most often ignored by the scenarios.
Relying on ICT in the future requires looking at energy and resource supply, which may be hindered by geopolitical tensions, collaborations between countries, and changes in governance. These tensions are discussed in French national studies (e.g. RTE and negaWatt) and in the narrative study Arup, but do not seem to influence how scenarios integrate ICT. The Extinction express scenario of the Arup study describes resource colonies on the Moon and in the deep sea.
To the best of our knowledge, however, no study mentions the geopolitical aspects related to the location of data centers and the installation of ICT infrastructures (e.g., underwater cables or satellites).
@@
Surprisingly, many disruptive technologies that have recently received attention and funding are not mentioned in any of the studies, which only build upon existing technologies. For example, few studies discuss 6G, and no studies discuss quantum informatics, intermittent computing, or DNA data storage. It is thus difficult to imagine how these innovations in ICT could be integrated into the scenarios.
An example of a study where the disruptive technologies of cyborg botany and olfactory notifications are considered in scenarios was recently presented in 2022 by French designers444 https://b-com.com/actualite/immersion-dans-le-numerique-responsable-de-2040. However, this work was not included in our analysis because of a lack of detail in the methodology.
Significant changes in the use of technology, such as using low-tech solutions or studying collapse informatics (Tomlinson et al., 2013), are also mostly absent from scenarios.
@@
Similar to the challenges in new technologies, very few general scenarios are peripheral, i.e., include unlikely and extreme events and consider a discontinuous path to the future. Typically, these disruptive scenarios are narratives.
In general scenarios, future ICT is quite similar to that of today. Our relationship to digital technologies is not deeply questioned, which may be why ICT is present in all scenarios without breakthrough innovations.
It is surprising in the context of Anthropocene that none of the studies analysed here propose disruptive scenarios where ICT would be less present or central in the society. The LIMITS community has nevertheless been active to challenge business-as-usual scenarios in the context of scarcity for some years now. The reasons of this non-disruption of ICT in the scenarios would need to be more deeply explored.
@@
The studies were all made from a climate change perspective, omitting biodiversity loss and other planetary boundaries, and focusing only on human futures without considering other species. These environmental problems may have other consequences on future scenarios. One that is directly related to ICT is water scarcity. Indeed, the manufacturing of ICT equipment and data center cooling have a substantial ecological footprint in terms of water use. Prospective studies on ICT should, therefore, broaden their narrations.
@@
Most innovations in the ICT sector and major industries of the domain come from Global North countries. As a consequence, the studies we have chosen were made in OECD countries, which clearly has an influence on scenarios. Although some studies (e.g., Arup) have been designed by people from several countries worldwide, our work is most likely not representative of visions from Global South studies, and future work should consider how to imagine ICT for the future in Global South countries.
6. Conclusion
In this paper, we have analyzed the place of ICT in prospective scenarios. We have highlighted some variables that influence the development of digital technologies and the areas of applications where ICT might be essential in the future. Our analysis demonstrated that the feasibility and consequences of the increasing development of digital technologies are not sufficiently considered. Through our discussion, we have shown that the studies hardly question humanity’s relationship to digital technologies, or the applications of ICT in human societies situated in a world of limits. The reason is probably that ICT is mostly seen as improving the society, as a solution to environmental challenges, or as immaterial. This shortcoming limits the possibility of imagining more diverse plausible futures for the ICT sector.
There is therefore still a need to design prospective general and non-general studies that could offer more diverse and systemic views of ICT, and that could highlight which undone science (Hess, 2015) should receive more attention. Prospective studies with a sustainability perspective for ICT can i) offer a more diverse and systemic view of the future of digital technologies, ii) highlight the need for discussing, structuring, or funding. As an example, (Deron and McDonald, 2022) demonstrate that, based on prospective co-design workshops, it is possible to find milestones in order to reach a future that is desirable of digital technologies. The milestones include among others sobriety measures, awareness, repairability and changes in some economic models, even if there is still work to be done to imagine the consequences on the others sectors of activities. They show samples of undone sciences (Hess, 2015) for which social and industrial movements, funding, and research are needed for a desirable and sustainable ICT. Design fiction has been promoted for many years now by the LIMITS community (e.g. (Tanenbaum et al., 2016)) for its ethical, rhetorical, and playful role. It is a tool to play with utopias and dystopias, to open dialogue between scientific researchers and society, and to think about the interplay between technological and sociological futures. The design of further studies must also be accompanied with a better understanding of the gaps in public knowledge and the misconceptions about sustainability in order to be able to impact the narrative around collapse (Tanenbaum et al., 2016). From our analysis, we additionally advocate for researching methodologies to propose more disruptive, and desirable for all, scenarios with a systemic point of view.
Appendix A Summary of studies
Table 2. Summary of studies reviewed for this paper
Study Summary & Scenarios
IPCC 2022 Report from IPCC Group III. This document does not correspond to a unique scenario, but describes the mitigation options.